OpenStudy (anonymous):

for what positive value of a does a-2sqrt(a+7) equals -4

OpenStudy (sidsiddhartha):

write it as a equation \[a-2\sqrt{a+7}=-4\\a+4=2\sqrt{a+7}\\now~ Square~ both~sides\\(a+4)^2=[2\sqrt{a+7}]^2\\a^2+8a+16=4(a+7)\\a^2+8a+16-4a-28=0\\a^2+4a-12=0\\now ~factor ~\it~and~take~positive~value~of~a\]

OpenStudy (sidsiddhartha):

can u solve it or need some help?

OpenStudy (anonymous):

yeh i can solve it now thanku:) @sidsiddhartha